# GASP: Gradient-Aware Shortest Path Algorithm for Boundary-Confined 2-Manifold Reeb Graph Visualization

### v1

@@ Abstract
Reeb graphs are an important tool for abstracting and representing the topological structure of a function defined on a manifold. We have identified three properties for faithfully representing Reeb graphs in a visualization. Namely, they should be constrained to the boundary, compact, and aligned with the function gradient. Existing algorithms for drawing Reeb graphs are agnostic to or violate these properties. In this paper, we introduce an algorithm to generate Reeb graph visualizations, called GASP, that is cognizant of these properties, thereby producing visualizations that are more representative of the underlying data. To demonstrate the improvements, the resulting Reeb graphs are evaluated both qualitatively and quantitatively against the geometric barycenter algorithm, using its implementation available in the Topology ToolKit (TTK), a widely adopted tool for calculating and visualizing Reeb graphs.
@@
Examples showing that the GASP algorithm (bottom) better accommodates the three desired properties for visualizing Reeb graphs, as compared to the geometric barycenter (GB) method, implemented in the Topology ToolKit (top). The complete Reeb graphs include both red and blue arcs, but red arcs highlight improvements over GB. (a) GASP constrains the path of the Reeb graph arcs such that they remain either inside or on the surface of the object, while many arcs generated by GB do not conform to the boundary of the model. (b) GASP chooses the shortest path between two critical points (while constrained by the boundary), while some GB arcs are unnecessarily long due to the desire for smoothness. (c) GASP maintains consistency with the gradient of the function, while in part due to smoothing, GB arcs travel nearly perpendicular to the direction of the function (left-to-right in this case).
@@ 1 Introduction
Since the Reeb graph is drawn based on a scalar function defined over a manifold, it should closely reflect those objects. To ensure this, we identified three essential properties and one optional property that a Reeb graph visualization should possess. First, because the scalar field is defined on the manifold, the arcs of the Reeb graph should remain confined within the manifold. Any arc extending beyond it would imply the Reeb graph exists outside the manifold (see LABEL:fig:teaser:outside top). Second, the Reeb graph arcs should be as short as possible to minimize visual clutter and to facilitate easier interpretation (see LABEL:fig:teaser:length top). Third, the arc orientations should closely align with the direction of the associated scalar field gradient. A mismatch between the gradient direction and the Reeb graph arcs can be confusing and make it harder to understand the scalar field (see LABEL:fig:teaser:direction top). Fourth, smoother arcs provide a clearer presentation of the paths, making them easier to track, but this property is optional, as it should not be at the cost of the other properties.
Several algorithms for drawing Reeb graphs and contour trees (loop-free Reeb graphs) have been proposed (see Section˜2.4). However, none explicitly identify or address all of these desired properties. Prior work on Reeb graph visualization has either focused on planar drawings or included only brief mentions within the broader context of Reeb graph construction [46, 18]. To the best of our knowledge, this paper is the first to focus exclusively on visualizing Reeb graphs in 3D embeddings.
@@ 2 Background on Reeb Graphs
We briefly present the technical definitions of isocontours and critical points that are building blocks of a Reeb graph construction. We then describe Reeb graph construction, computation, and visualization.
@@ 2.1 Isocontours and Critical Points
Isocontours are a foundational tool in visualization. Isocontours represent a locus of points in a 2-manifold that attain a fixed function value, known as an isovalue. We restrict our discussion to scalar functions sampled on a $2$-dimensional manifold, as other types of functions and manifolds are beyond the scope of this paper. Mathematically, if $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$ is a scalar function defined on an $2$-dimensional manifold, $\mathcal{M}$}, then the isocontour $\mathcal{L}$ for isovalue $k$ can be represented as $\mathcal{L}(k)\equiv\{P:P\in{\mathcal{M}}\land f(P)=k\}$, where $P$ denotes domain positions with function value $k$. The change in data values over the domain can be understood by visualizing the evolution of isocontours for isovalues spaced at regular intervals. LABEL:fig:reebGraphConcept:a visualizes isocontours (in orange) of a height function, $f$, for various isovalues, $f_{i}$, mapped on a torus. A single isovalue may produce a single or multiple connected components. As observed in LABEL:fig:reebGraphConcept:a, the isocontours immediately beyond split and merge saddles comprise two connected components.
@@ 2.2 The Reeb Graph
The Reeb graph is a structural abstraction that provides insight into the topological skeleton of scalar field data. Formally, the Reeb graph tracks the connected components of isocontours in the domain as $f$ is swept from $\infty\rightarrow+\infty$ [20]. The Reeb graph of a Morse function, $f$, defined on a simply connected manifold, $\mathcal{M}$}, is loop-free and is referred to as a contour tree [10]. As shown in LABEL:fig:reebGraphConcept:b, each connected component of isocontour (orange ellipses) can be contracted to a single point (depicted as blue dots). Connecting the points representing connected components forms arcs or edges ($E$) of a Reeb graph, as portrayed in LABEL:fig:reebGraphConcept:c. In summary, the arcs ($E$) of a Reeb graph indicate connected components in the domain with the same topology, and critical points ($C$) in the Reeb graph represent nodes that indicate topological changes. Reeb graphs, therefore, provide a compact and abstract visual representation of the data topology.
@@ 2.3 Reeb Graph Computation
Considering the high utility of Reeb graphs in the analysis of complex data, several prior works focused on methods for efficient computation and visualization of Reeb graphs. Shinagawa and Kunii [50] proposed $O(n^{2})$ algorithm for Reeb graph construction, where $n$ represents the number of vertices in a 2-manifold triangular mesh. Since then, multiple new algorithms were proposed that increased the efficiency of Reeb graph constructions [13, 16, 17, 56, 47]. Harvey and Wang [29] achieved an optimal expected running time of $O(m\log m)$ through a randomized dynamic-connectivity strategy, and Parsa [44] derandomized this approach to obtain the same bound in the worst case. Gueunet et al. [26] presented a task-based parallel algorithm for Reeb Graph calculation and visualization, leveraging merge and split trees constructed from scalar field data. It uses Fibonacci heaps to efficiently handle dynamic operations like merging and splitting components. TTK implements this approach for Reeb graph computation. While this work targets parallel performance, its sequential core follows Parsa’s framework, which itself builds on the randomized dynamic-connectivity formulation introduced by Harvey and Wang.
@@ 2.4 Reeb Graph Visualization
Several approaches have been used to visualize Reeb graphs, contour trees, and general graphs. Strodthoff and Jüttler [53] devised a layered-Reeb graph construction algorithm for efficient computation and visualization of Reeb graphs for solids with boundary embedded in 3D. Shinagawa et al. [51] devised a 2D layout for the representation of Reeb graphs. Pascucci et al. [45] devised “toporrery” as a way to draw contour trees in a radial manner that avoids the self-intersection of edges in planar methods. Heine et al. [30] leveraged advances in graph drawing techniques [24, 4] for drawing contour trees in 2D layouts. 3D medial axis algorithms [14] deal with a related problem, in that they produce a skeleton of a manifold. However, they do not have the constraints introduced by the scalar function defined on the manifold.
@@
Refer to captionRefer to captionRefer to caption
Figure 2: High-level algorithm description. (a) A Reeb graph is calculated. Although it is shown on the mesh, the calculation only provides a graph object with critical points as nodes and arcs as edges. (b) Our approach first decomposes the mesh into a set of topological cylinders, one per edge. (c) A series of isocontours are calculated for each topological cylinder, and candidate points for the arc path are calculated by (d) the boundary approach using the contours themselves or (e) the interior approach by finding points inside the contours. (f) Once the candidate points are found, graph edges are formed between the critical points and the adjacent contours and between adjacent contours, and the output arc is calculated as the shortest path between the critical points. (g) Finally, the arcs from all topological cylinders are assembled into a final Reeb graph.
@@
The Topology ToolKit (TTK) [55] is a frequently used tool for visualizing Reeb graphs, which uses the geometrical barycenter algorithm [46]. TTK’s method for visualizing Reeb graphs involves three steps [54, 46]. First, Reeb graph edges connect critical points directly via straight lines (see LABEL:fig:TTK-Bird-lines). Arcs are then produced using an arc sampling parameter that evenly divides the straight edge into sub-components (see LABEL:fig:TTK-Bird-arcs). Then, each sample is placed at the geometrical barycenter of the isocontours, and, finally, the arcs are smoothed geometrically to improve the aesthetics of the representation (see LABEL:fig:TTK-Bird-Curves). TTK produces smooth, aesthetically pleasing Reeb graph arcs; however, in many cases, it fails to produce Reeb graph arcs with the three essential properties we identified. First, the process used by TTK often results in arcs that extend beyond the model’s boundary (see the red arcs in LABEL:fig:teaser:outside top). Further, the sampling and smoothing processes often lead to unnecessarily long paths, which cause visual clutter (see the red arcs in LABEL:fig:teaser:length top). Finally, the process of smoothing can introduce significant deviations from the function gradient direction (see the red arcs in LABEL:fig:teaser:direction top).
@@ 3 Literature Review
3.1 Reeb Graphs for Data Analysis
@@
Reeb graphs play a pivotal role in data analysis. Reeb graphs and contour trees (as well as other topology-based techniques) are often used as a way to compare and quantitatively measure the similarity of features across complex scalar fields, as surveyed in [59]. Various metrics, including interleaving [38], edit [52], and stable [8] distance have been proposed to quantitatively measure the distance between merge trees that are foundational to construction of Reeb graphs and contour trees. Recently, functional distortion metric [5] and intrinsic interleaving distances between merge trees [35] were used to compare Reeb graphs quantitatively. Chen et al. [11] analyzed time-dependent multi-fluid data by studying Reeb graphs of fluid density distributions. Weber et al. [58] proposed feature tracking using Reeb graphs in time-varying data for combustion applications. Reeb graphs are also a fundamental tool in shape analysis in computer graphics and other domains [7, 6, 43]. Natali et al. [40] proposed utilizing Reeb graphs for extracting the topological skeleton of point cloud data. Given the widespread use of Reeb graphs in data analysis, accurate visualization methods for Reeb graphs are a critical component of that analysis.
3.2 Topological Visualizations
@@
Topological visualization is a powerful tool used to concisely convey the scale and position of important data features to facilitate analysis of complex scientific data [9, 48, 41, 42]. While isocontour [36], critical point [39], and Morse-Smale complex [21] techniques explicitly convey position of important data features, persistence diagram and persistence curve techniques [12] explicitly convey scale of data features. Reeb graphs [20] and contour trees [10] provide topological abstractions that convey feature positions through critical points and scale through arc length in a planar representation. The discrete graph-based representation of these topological visualization types, however, can make their perception and data analysis tasks difficult. Topological simplification [22, 27, 18] is, therefore, a standard method used to reduce the level of detail or noise in data to make topological visualizations more interpretable. Recent work by Athawale et al. [3] and previous work on vector field topology [34, 23] showed a limited capability of users to perceive topological features conveyed by various methods. Their study concluded a need to improve existing visualization designs to make topological visualizations more perceptible. In this paper, we propose a novel algorithm, GASP, to enhance the faithfulness of Reeb graph visualization compared to the geometric barycenter algorithm.
@@ 4 GASP (Gradient-Aware Shortest Path) Reeb Graphs
We target generating Reeb graphs that are better representations of the model and function used to generate them. To do this, we begin with a high-level conceptual description. The input to our approach is a 2-manifold embedded in 3D space, $\mathcal{M}$}, a Morse function, $f$, and the Reeb graph consisting of a set of critical points, $C$, and Reeb graph edges, $E$, which connect pairs of critical points (i.e., $E_{i}=(C_{j},C_{k})$). Our process, shown in Fig.˜2, contains three main steps.
@@ 4.2.2 Candidate Points
The next step involves generating a series of candidate points, $p$, using the contours from the prior step, which will be used to identify the Reeb graph path. We present two variations. The first variation called the boundary approach (see LABEL:fig:conceptual:boundary), selects a series of candidate points along the contour boundary itself. As will be noted in the implementation, we directly use the points produced by isocontouring. The second variation called the interior approach (see LABEL:fig:conceptual:interior), generates a series of candidate points in a grid-like pattern inside the contour and some user-set distance from the boundary (denoted by the orange stripes).
@@ 4.3 Final Assembly
Figure 4: (a) Torus (left) and Modified Torus (right) are (b) binned by function value and have a rough cut in cyan and blue performed between the critical points in red. Then, the triangles in blue have the (c) source cut applied, (d) followed by the destination cut. (e-g) Examples of different candidate connected component validation criteria after a cut, where valid cylinders are blue, and invalid ones are gray. (e) After a cut, a valid connected component will connect the two critical points ($c_{j}$ and $c_{k}$), e.g., $\overline{{{\mathbb{C}}}}_{2}$, whereas $\overline{{{\mathbb{C}}}}_{1}$ and $\overline{{{\mathbb{C}}}}_{3}$ only connect to one critical point, $c_{j}$ and $c_{k}$, respectively. (f) After a cut, a valid connected component will not contain any other critical points, e.g., $\overline{{{\mathbb{C}}}}_{2}$, where $\overline{{{\mathbb{C}}}}_{1}$ contains the extra critical point $c_{l}$. (g) Special care must be taken for Reeb graph edges with duplicity as multiple valid topological cylinders will be identified (e.g., both $\overline{{{\mathbb{C}}}}_{1}$ and $\overline{{{\mathbb{C}}}}_{2}$) and must be matched to those edges.

### v2

@@ Abstract
Reeb graphs are an important tool for abstracting and representing the topological structure of a function defined on a manifold. We have identified three properties for faithfully representing Reeb graphs in a visualization: they should be constrained to the boundary, compact, and aligned with the function gradient. Existing algorithms for drawing Reeb graphs are agnostic to or violate these properties. In this paper, we introduce an algorithm to generate Reeb graph visualizations, called GASP, that is cognizant of these properties, thereby producing visualizations that are more representative of the underlying data. To demonstrate the improvements, the resulting Reeb graphs are evaluated both qualitatively and quantitatively against the geometric barycenter algorithm, using its implementation available in the Topology ToolKit (TTK), a widely adopted tool for calculating and visualizing Reeb graphs.
@@
Examples showing that the GASP algorithm (bottom) better accommodates the three desired properties for visualizing Reeb graphs, as compared to the geometric barycenter (GB) method, implemented in the Topology ToolKit (top). The complete Reeb graphs include both red and blue arcs, but red arcs highlight improvements over GB. (a) GASP constrains the path of the Reeb graph arcs such that they remain either inside or on the surface of the object, whereas many arcs generated by GB do not conform to the boundary of the model. (b) GASP chooses the shortest path between two critical points (while constrained by the boundary), but some GB arcs are unnecessarily long due to the desire for smoothness. (c) GASP maintains consistency with the gradient of the function, while in part due to smoothing, GB arcs travel nearly perpendicular to the direction of the function (left-to-right in this case).
@@ 1 Introduction
Since the Reeb graph is drawn based on a scalar function defined over a manifold, it should closely reflect those objects. To ensure this, we identified three essential properties and one optional property that a Reeb graph visualization should possess. First, because the scalar field is defined on the manifold, the arcs of the Reeb graph should remain confined within the manifold. Any arc extending beyond it would imply the Reeb graph exists outside the manifold (see LABEL:fig:teaser:outside top). Second, the Reeb graph arcs should be as short as possible to minimize visual clutter and to facilitate easier interpretation (see LABEL:fig:teaser:length top). Third, the arc orientations should closely align with the direction of the associated scalar field gradient. A mismatch between the gradient direction and the Reeb graph arcs can be confusing and make it harder to understand the scalar field (see LABEL:fig:teaser:direction top). Fourth, smoother arcs provide a clearer presentation of the paths, making them easier to track, but this property is optional, as it should not be at the cost of the other properties. This manuscript has been authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The publisher, by accepting the article for publication, acknowledges that the U.S. Government retains a non-exclusive, paid up, irrevocable, world-wide license to publish or reproduce the published form of the manuscript, or allow others to do so, for U.S. Government purposes. The DOE will provide public access to these results in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
Several algorithms for drawing Reeb graphs and contour trees (loop-free Reeb graphs) have been proposed (see Section˜2.4). However, none explicitly identify or address all of these properties. Prior work on Reeb graph visualization has focused on planar drawings or included only brief mentions within the context of Reeb graph construction [46, 18]. To the best of our knowledge, this paper is the first to focus exclusively on visualizing $2$-manifold Reeb graphs in 3D embeddings.
@@ 2 Background on Reeb Graphs
We briefly present the technical definitions of isocontours and critical points that are the building blocks of a Reeb graph construction. We then describe Reeb graph construction, computation, and visualization.
@@ 2.1 Isocontours and Critical Points
Isocontours are a foundational tool in visualization. Isocontours represent a locus of points in a $2$-dimensional manifold that attain a fixed function value, known as an isovalue. We restrict our discussion to scalar functions sampled on a $2$-manifold, as other types of functions and manifolds are beyond the scope of this paper. Mathematically, if $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$ is a scalar function defined on a $2$-manifold, $\mathcal{M}$}, then the isocontour $\mathcal{L}$ for isovalue $k$ can be represented as $\mathcal{L}(k)\equiv\{P:P\in{\mathcal{M}}\land f(P)=k\}$, where $P$ denotes domain positions with function value $k$. The change in data values over the domain can be understood by visualizing the evolution of isocontours for isovalues spaced at regular intervals. LABEL:fig:reebGraphConcept:a visualizes isocontours (in orange) of a height function, $f$, for various isovalues, $f_{i}$, mapped on a torus. A single isovalue may produce a single or multiple connected components. As observed in LABEL:fig:reebGraphConcept:a, the isocontours immediately beyond split and merge saddles comprise two connected components.
@@ 2.2 The Reeb Graph
The Reeb graph is a structural abstraction that provides insight into the topological skeleton of scalar field data. Formally, the Reeb graph tracks the connected components of isocontours in the domain as $f$ is swept from $\infty\rightarrow+\infty$ [20]. On a simply connected manifold, $\mathcal{M}$}, the Reeb graph of a Morse function, $f$ is loop-free (a generalization of contour tree)  [10]. As shown in LABEL:fig:reebGraphConcept:b, each connected component of an isocontour (orange ellipses) can be contracted to a single point (depicted as blue dots). Connecting the points representing connected components forms arcs or edges ($E$) of a Reeb graph, as portrayed in LABEL:fig:reebGraphConcept:c. In summary, the arcs ($E$) of a Reeb graph indicate connected components in the domain with the same topology, and critical points ($C$) in the Reeb graph represent nodes that indicate topological changes. Reeb graphs, therefore, provide a compact and abstract visual representation of the data topology.
@@ 2.3 Reeb Graph Computation
Considering the high utility of Reeb graphs in the analysis of complex data, several prior works focused on methods for efficient computation and visualization of Reeb graphs. Shinagawa and Kunii [51] proposed $O(n^{2})$ algorithm for Reeb graph construction, where $n$ represents the number of vertices in a $2$-manifold triangular mesh. Since then, multiple new algorithms were proposed that increased the efficiency of Reeb graph constructions [13, 16, 17, 57, 47]. Harvey and Wang [29] achieved an optimal expected running time of $O(m\log m)$ through a randomized dynamic-connectivity strategy, and Parsa [44] derandomized this approach to obtain the same bound in the worst case. Gueunet et al. [26] presented a task-based parallel algorithm for Reeb graph calculation and visualization, leveraging merge and split trees constructed from scalar field data. It uses Fibonacci heaps to efficiently handle dynamic operations like merging and splitting components. TTK implements this approach for Reeb graph computation. Although this work targets parallel performance, its sequential core follows Parsa’s framework, which itself builds on the randomized dynamic-connectivity formulation introduced by Harvey and Wang.
@@ 2.4 Reeb Graph Visualization
Several approaches have been used to visualize Reeb graphs, contour trees, and general graphs. Strodthoff and Jüttler [54] devised a layered-Reeb graph construction algorithm for efficient computation and visualization of Reeb graphs for solids with boundary embedded in 3D. Shinagawa et al. [52] devised a 2D layout for the representation of Reeb graphs. Pascucci et al. [45] devised “toporrery” as a way to draw contour trees in a radial manner that avoids the self-intersection of edges in planar methods. Heine et al. [30] leveraged advances in graph drawing techniques [24, 4] for drawing contour trees in 2D layouts. 3D medial axis algorithms [14] deal with a related problem, in that they produce a skeleton of a manifold. However, they do not have the constraints introduced by the scalar function defined on the manifold. Rehal and Sen [48] proposed computing all non-homotopic paths between points on a surface or in its ambient space using routing graphs derived from handle and tunnel loops, offering a topology-aware approach conceptually related to skeleton and graph generation methods.
@@
The Topology ToolKit (TTK) [56] is a frequently used tool for visualizing Reeb graphs that uses the geometrical barycenter algorithm [46]. TTK’s method for visualizing Reeb graphs involves three steps [55, 46]. First, Reeb graph edges connect critical points directly via straight lines (see LABEL:fig:TTK-Bird-lines). Arcs are then produced using an arc sampling parameter that evenly divides the straight edge into sub-components (see LABEL:fig:TTK-Bird-arcs). Then, each sample is placed at the geometrical barycenter of the isocontours, and, finally, the arcs are smoothed geometrically to improve the aesthetics of the representation (see LABEL:fig:TTK-Bird-Curves). TTK produces smooth, aesthetically pleasing Reeb graph arcs; however, in many cases, it fails to produce Reeb graph arcs with the three essential properties we identified. First, the process used by TTK often results in arcs that extend beyond the model’s boundary (see the red arcs in LABEL:fig:teaser:outside top). Further, the sampling and smoothing processes often lead to unnecessarily long paths, which cause visual clutter (see the red arcs in LABEL:fig:teaser:length top). Finally, the process of smoothing can introduce significant deviations from the function gradient direction (see the red arcs in LABEL:fig:teaser:direction top).
@@ 3 Literature Review
Reeb Graphs for Data Analysis
@@
Reeb graphs play a pivotal role in data analysis. Reeb graphs and contour trees (as well as other topology-based techniques) are used as a way to compare and measure the similarity of features across complex scalar fields, as surveyed in [60]. Various metrics, including interleaving [38], edit [53], and stable [8] distance, have been proposed to measure the distance between merge trees that are foundational to the construction of Reeb graphs and contour trees. Recently, functional distortion metric [5] and intrinsic interleaving distances between merge trees [35] were used to compare Reeb graphs quantitatively. Chen et al. [11] analyzed time-dependent multi-fluid data by studying Reeb graphs of fluid density distributions. Weber et al. [59] proposed feature tracking using Reeb graphs in time-varying data for combustion applications. Reeb graphs are also a fundamental tool in shape analysis in computer graphics and other domains [7, 6, 43]. Natali et al. [40] proposed utilizing Reeb graphs for extracting the topological skeleton of point cloud data. Given the widespread use of Reeb graphs in data analysis, accurate visualization methods for Reeb graphs are a critical component of that analysis.
Topological Visualizations
@@
Topological visualization is a powerful tool used to concisely convey the scale and position of important data features to facilitate analysis of complex scientific data [9, 49, 41, 42]. While isocontour [36], critical point [39], and Morse-Smale complex [21] techniques explicitly convey the position of important data features, persistence diagram and persistence curve techniques [12] convey the scale of data features. Reeb graphs [20] and contour trees [10] provide topological abstractions that convey feature positions through critical points and scale through arc length in a planar representation. The discrete graph-based representation of these topological visualization types, however, can make their perception and data analysis tasks difficult. Topological simplification [22, 27, 18] is, therefore, a standard method used to reduce the level of detail or noise in data to make topological visualizations more interpretable. Recent work by Athawale et al. [3] and previous work on vector field topology [34, 23] showed a limited capability of users to perceive topological features conveyed by various methods. The studies concluded a need to improve existing visualization designs to make topological visualizations more perceptible. In this paper, we propose a novel algorithm, GASP, to enhance the faithfulness of Reeb graph visualization compared to the geometric barycenter algorithm.
Refer to captionRefer to captionRefer to caption
Figure 3: GASP algorithm description. (a) A Reeb graph is calculated. While shown on the mesh, the calculation provides only a graph object with critical points as nodes and arcs as edges. (b) GASP first decomposes the mesh into a set of topological cylinders, one per edge. (c) A series of isocontours is calculated for each topological cylinder, and candidate points for the arc path are calculated by (d) the boundary approach using the contours themselves or (e) the interior approach by finding points inside the contours. (f) Once the candidate points are found, graph edges are formed between the critical points and the adjacent contours and between adjacent contours, and the output arc is calculated as the shortest path between the critical points. (g) Finally, the arcs from all topological cylinders are assembled into a final Reeb graph.
@@ 4 GASP (Gradient-Aware Shortest Path) Reeb Graphs
We target generating Reeb graphs that are better representations of the model and function used to generate them. To do this, we begin with a high-level conceptual description. The input to our approach is a 2-manifold embedded in 3D space, $\mathcal{M}$}, a Morse function, $f$, and the Reeb graph consisting of a set of critical points, $C$, and Reeb graph edges, $E$, which connect pairs of critical points (i.e., $E_{i}=(C_{j},C_{k})$). Our process, shown in Fig.˜3, contains three main steps.
@@ 4.2.2 Candidate Points
The next step involves generating a series of candidate points, $p$, using the contours from the prior step, which will be used to identify the Reeb graph path. We present two variations. The first variation, called the boundary approach (see LABEL:fig:conceptual:boundary), selects a series of candidate points along the contour boundary itself. As will be noted in the implementation, we directly use the points produced by isocontouring. The second variation, called the interior approach (see LABEL:fig:conceptual:interior), generates a series of candidate points in a grid-like pattern inside the contour and some user-set distance from the boundary (denoted by the orange stripes).
@@ 4.3 Final Assembly
Figure 4: (a) Torus (left) and modified torus (right) are (b) binned by function value and have a rough cut in cyan and blue performed between the critical points in red. Then, the triangles in blue have the (c) source cut applied, (d) followed by the destination cut. (e-g) Examples of different candidate connected component validation criteria after a cut, where valid cylinders are blue, and invalid ones are gray. (e) After a cut, a valid connected component will connect the two critical points ($c_{j}$ and $c_{k}$), e.g., $\overline{{{\mathbb{C}}}}_{2}$, whereas $\overline{{{\mathbb{C}}}}_{1}$ and $\overline{{{\mathbb{C}}}}_{3}$ connect to only one critical point, $c_{j}$ and $c_{k}$, respectively. (f) After a cut, a valid connected component will not contain any other critical points, e.g., $\overline{{{\mathbb{C}}}}_{2}$, where $\overline{{{\mathbb{C}}}}_{1}$ contains the extra critical point $c_{l}$. (g) Special care must be taken for Reeb graph edges with duplicity as multiple valid topological cylinders will be identified (e.g., both $\overline{{{\mathbb{C}}}}_{1}$ and $\overline{{{\mathbb{C}}}}_{2}$) and must be matched to those edges.